OpenStudy (anonymous):

please help me im takeing a test and i dont know the answer Find the measure of the angle. Angles A and B are supplementary. m∠A = 24° m∠B = A. 66° B. 76° C. 156° D. 166°

OpenStudy (anonymous):

C . 156° Just x + 24 = 180 x = 180 - 24 x = 156 :)

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

no problem :)